OpenStudy (anonymous):

will medal The measurement of one angle of a right triangle is 61������. What is the measurement of the missing angle?

OpenStudy (anonymous):

@Data_LG2 @Destinymasha @Here_to_Help15

OpenStudy (anonymous):

total angle of a triangle is 180 degrees. you are given: 61 degrees and 90 degrees since it is a right triangle. so 180= 61+ 90 + x, let x the missing angle solving for x.. x=180-90-61 x=?

OpenStudy (anonymous):

X= 61 that's what i got

OpenStudy (anonymous):

wait a min no its not that let me check it on my calculator

OpenStudy (anonymous):

X= 29

OpenStudy (anonymous):

So @Data_LG2 is right and i am right

OpenStudy (anonymous):

yes, x=29 sorry late response :)

OpenStudy (anonymous):

ok it all good